At which point on the following curve does the tangent line has the largest slope?


𝑦 = 1 + 40π‘₯^3 βˆ’ 3π‘₯^5


1
Expert's answer
2022-03-31T02:56:39-0400

"y'\left( x \right) \rightarrow \max \\y'=120x^2-15x^4\rightarrow \max \\{x^2}_{\max}=\frac{-120}{2\cdot \left( -15 \right)}=4\\x=\pm 2"